A manager has been using a certain technique to forecast demand for gallons of ice cream for the past six periods. Actual and predicted amounts are shown below. Calculate MAD, MSE, MAPE

    

Period

Demand

Forecast

1

90

87

2

85

88

3

91

87

4

92

89

5

85

90

6

88

92
Answer #1

MAD(Mean Absolute deviation) = Sum of Absolute deviation/Number of observations

= (3 + 3 + 4 + 3 + 5 + 4)/6

= 22/6

= 3.67

MSE (Mean Square Error) = Sum of Square Errors/Number of observations

= (9 + 9 + 16 + 9 + 25 + 16)/6

= 14

MAPE(Mean Absolute Percentage Error) = Sum of percentage errors/Number or observations

= (3.33 + 3.53 + 4.40 + 3.26 + 5.88 + 4.55)/6

= 4.16%